OpenStudy (anonymous):

Plastic storage bins are designed so they can be nested. Four 10-gallon bins stacked together measure 26 inches high, and 9 stacked bins measure 41 inches high. Write an equation in the form y = mx + b to describe the relationship between x, the number of nested bins, and y, the height of the stack.

OpenStudy (danjs):

they give you two points of data here. Let x = Number of bins y = Height of the stack Points (x,y) given (4,26) and (9, 41) calculate the slope first

OpenStudy (danjs):

m = slope (x1,y1) = (4,26) (any point on the line) Point-Slope Form of the line y - y1 = m*(x-x1) Put in data, and rearrange to y =m*x+b form.